# On a discrete version of the position-momentum commutation relation

### v4

@@ V Discrete version of the position-momentum commutation relation
In the continuous case, the usual position and momentum satisfy the remarkable relation
@@
Table 2: Case d=11d\!=\!11: $\varphi_{1},\,\varphi_{2},...,\varphi_{d}$ are eigenstates of $\mathfrak{F}$.
$k$ $\parallel\mathfrak{F}\varphi_{k}\!-\!\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!+\!\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!-\!\mathrm{i}\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!+\!\mathrm{i}\varphi_{k}\parallel$
@@
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}$ (25)
One can directly see that this relation is not satisfied, but by computing the eigenvalues of the operator
$\begin{array}[]{l}\hat{\mathfrak{C}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\quad\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\end{array}$ (26)
with the matrix elements
$\begin{array}[]{l}\langle n|\hat{\mathfrak{C}}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k(n-m)]-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (27)
for different values of $d$, one arrives at the conclusion that the eigenvalues $\lambda$ with $|\lambda|<0.001$ represent $36\%$ in the case d=11d\!=\!11, represent $64\%$ in the case d=31d\!=\!31, represent $77\%$ in the case d=61d\!=\!61 and $84\%$ in the case d=101d\!=\!101 (see Table I). So, for $d$ large enough, a significant part of the eigenvalues of $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ are almost null.
For any $\varepsilon\!>\!0$, there exists $d_{\varepsilon}$ such that the set
@@
$\begin{array}[]{l}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\frac{d}{2\pi}\psi\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.\end{array}$ (29)
@@
that is $\psi\!\in\!\mathcal{S}_{\varepsilon\sqrt{k}}$.
@@
By using the relation
$\begin{array}[]{r}(\hat{\mathfrak{F}}\hat{\mathfrak{q}}\hat{\mathfrak{F}}^{\dagger})\psi(n)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}m(k-n)]\,m\,\psi(k)\\ =\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{d}m(k-n)]\,(-m)\,\psi(k)\\ =\!-(\hat{\mathfrak{F}}^{\dagger}\hat{\mathfrak{q}}\hat{\mathfrak{F}})\psi(n)\!=\!-\hat{\mathfrak{p}}\psi(n)\end{array}$
we get $\hat{\mathfrak{F}}\hat{\mathfrak{q}}\hat{\mathfrak{F}}^{\dagger}\!=\!-\hat{\mathfrak{p}}$ and
[…]
that is
$\hat{\mathfrak{F}}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\hat{\mathfrak{F}}$. (31)
From the relation
[…]
$\psi\!\in\!\mathcal{S}_{\varepsilon}\quad\Rightarrow\quad\hat{\mathfrak{F}}\psi\!\in\!\mathcal{S}_{\varepsilon}.$ (33)
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{0}|\leq|\lambda_{1}|\leq...\leq|\lambda_{d-1}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
$\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\ \varphi_{\kappa}.$ (34)
The relation
$\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\,\varphi_{\kappa}\quad\Rightarrow\quad\hat{\mathfrak{C}}\hat{\mathfrak{F}}\varphi_{k}\!=\!\hat{\mathfrak{F}}\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\ \hat{\mathfrak{F}}\,\varphi_{\kappa}$
and the numerical data (see Table 2) suggest that the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ are also eigenstates of $\hat{\mathfrak{F}}$, namely
$\hat{\mathfrak{F}}\varphi_{n}\!=\!(-\mathrm{i})^{n}\ \varphi_{n}.$ (35)
Particularly, the discrete Fourier transform admits the spectral resolution
$\hat{\mathfrak{F}}\!=\!\sum_{n=0}^{d-1}(-\mathrm{i})^{n}\ |\varphi_{n}\rangle\langle\varphi_{n}|.$ (36)
VI A discrete version of the uncertainty relation
A direct consequence of (29) is
$|\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|\approx\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.$ (37)
$\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\frac{|\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi|}{2}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon},$ (38)
that is
$\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (39)
is approximately satisfied in $\mathcal{S}_{\varepsilon}$.
VII Discrete-variable Gaussian functions
Figure 1: Discrete Gaussian functions: $\mathbf{g}_{3}$ and $\mathbf{g}_{1/3}$ in the case d=31d\!=\!31.
$\begin{array}[]{l}g_{\kappa}(q)\!=\!\mathrm{Exp}[-\frac{\kappa\pi}{h}q^{2}],\end{array}$ (40)
satisfies the relation
$\hat{F}[g_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\,g_{1/\kappa}.$ (41)
The corresponding discrete-variable Gaussian function
$\begin{array}[]{l}\mathfrak{g}_{\kappa}\!:\!\{-s,\!-s\!+\!\!1,...,s\!-\!\!1,\!s\}\!\rightarrow\!\mathbb{R},\\[2.84526pt] \mathfrak{g}_{\kappa}(n)=\sum\limits_{\alpha=-\infty}^{\infty}\mathrm{Exp}[-\frac{\kappa\pi}{d}(n+\alpha d)^{2}],\end{array}$ (42)
satisfies the relation [2]
$\hat{\mathfrak{F}}[\mathfrak{g}_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\ \mathfrak{g}_{1/\kappa}$ (43)
similar to (41). It is a particular case of Jacobi $\theta$ function [2, 7]. Because $\hat{\mathfrak{F}}$ is a unitary transform, the corresponding normalized discrete Gaussian functions (see Figure 1)
[…]
satisfy the relation
$\hat{\mathfrak{F}}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}.$ (45)
For $d$ large enough and small $\delta>0$, there exists $\eta>0$ such that, for $\kappa\!\in\!(\frac{1}{\eta},\eta)$, the most significant coordinates of $\mathbf{g}_{\kappa}$ in the eigenbasis $\{\varphi_{1},\,\varphi_{2},\,...,\,\varphi_{d}\}$ of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ are those corresponding to the functions $\varphi_{k}$ with $|\lambda_{k}|<\delta$ (see Table 3). We can consider that
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \mathbf{g}_{\kappa}\ \mbox{with}\ \ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ (46)
(see Tables 3, 4 and Figure 2). From (35), we get
$|\langle\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\theta_{n}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\mathfrak{F}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\varphi_{n}|\mathbf{g}_{1/\kappa}\rangle|,$ (47)
in agreement with the numerical data from Table 3. Since $\hat{\mathfrak{F}}$ is a unitary transform, we get
@@ VII Discrete-variable Gaussian functions
$\parallel\mathfrak{C}\mathbf{g}_{\kappa}\parallel=\parallel\mathfrak{C}\mathbf{g}_{1/\kappa}\parallel,$ (49)
@@
$\langle\mathbf{g}_{1/\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{1/\kappa}\rangle\!=\!\langle\hat{\mathfrak{F}}\mathbf{g}_{\kappa}|\hat{\mathfrak{F}}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle\!=\!\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle,$ (50)
@@
Table 4: $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\mathbf{g}_{\kappa}\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\mathbf{g}_{\kappa}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\mathbf{g}_{\kappa}||$
@@
Table 5: $|\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle|\approx\frac{d}{2\pi}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $|\ |\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle|-\frac{d}{2\pi}\ |$
@@
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{and}\quad\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (51)
are approximately satisfied for any $\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}$ (See Figure 2 and Figure 3).
Figure 2: Relation $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}$ in Gaussian case : $\parallel([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi})\mathbf{g}_{k}||$ for d=11d\!=\!11 and d=31d\!=\!31.
Figure 3: Relation $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ in Gaussian case : $|\ |\langle\mathbf{g}_{k}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{k}\rangle|-\frac{d}{2\pi}\ |$ for d=11d\!=\!11 and d=31d\!=\!31.
@@
$\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}\quad\Rightarrow\quad\hat{\mathfrak{F}}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}\!\in\!\mathcal{G}_{\delta}.$ (52)
VIII Discrete creation operator
The operators
[…]
can be regarded as a discrete version of the creation and annihilation operators
[…]
used in the continuous variable case. For small $\varepsilon\!>\!0$ and any $\psi\!\in\!\mathcal{S}_{\varepsilon}$, we have
$\hat{\mathfrak{a}},\hat{\mathfrak{a}}^{\dagger}]\!\approx\!1.$ (55)

### v5

@@ III Discrete version of the position-momentum commutation relation
In the continuous case, the usual position and momentum satisfy the remarkable relation $\hat{q},\hat{p}]={\rm i}\hbar$, that is
@@
$\hat{\mathbf{q}},\hat{\mathbf{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}$ (39)
One can see that this relation is not satisfied in the whole $\mathcal{H}$, but by computing the eigenvalues of the operator $\mathbf{C}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\end{array}$ (40)
with the matrix elements
$\begin{array}[]{l}\langle n|\mathbf{C}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}k(n-m)}-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (41)
for different values of $d$, one arrives at the conclusion that the eigenvalues $\lambda$ with $|\lambda|<0.001$ represent $36\%$ in the case d=11d\!=\!11, represent $64\%$ in the case d=31d\!=\!31, represent $77\%$ in the case d=61d\!=\!61 and $84\%$ in the case d=101d\!=\!101 (see Table I). So, for $d$ large enough, a significant part of the eigenvalues of $\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ are almost null.
For any $\varepsilon\!>\!0$, there exists $d_{\varepsilon}$ such that the set
[…]
$\begin{array}[]{l}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx{\rm i}\frac{d}{2\pi}\psi\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.\end{array}$ (43)
[…]
that is $\psi\!\in\!\mathcal{S}_{\varepsilon\sqrt{k}}$.
III.1 A spectral decomposition of the Fourier transform
By using the relation
$(\mathbf{F}\hat{\mathbf{q}}\mathbf{F}^{\dagger})\psi(n)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{e}^{\frac{2\pi\mathrm{i}}{d}m(k-n)}\,m\,\psi(k)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{e}^{-\frac{2\pi\mathrm{i}}{d}m(k-n)}\,(-m)\,\psi(k)\!=\!-(\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F})\psi(n)\!=\!-\hat{\mathbf{p}}\psi(n)$
we get $\mathbf{F}\hat{\mathbf{q}}\mathbf{F}^{\dagger}\!=\!-\hat{\mathbf{p}}$ and
$\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\!=\!\mathbf{F}\hat{\mathbf{q}}\,\hat{\mathbf{p}}-\mathbf{F}\hat{\mathbf{p}}\,\hat{\mathbf{q}}\!=\!\mathbf{F}\hat{\mathbf{q}}\,\mathbf{F}^{\dagger}\hat{\mathbf{F}}-\mathbf{F}\mathbf{F}^{\dagger}\hat{\mathbf{q}}\mathbf{F}\,\hat{\mathbf{q}}\!=\!-\hat{\mathbf{p}}\,\hat{\mathbf{q}}\mathbf{F}-\hat{\mathbf{q}}\mathbf{F}\,\hat{\mathbf{q}}\mathbf{F}^{\dagger}\mathbf{F}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{F},$
that is
$\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{F}.$ (45)
From the relation
[…]
$\psi\!\in\!\mathcal{S}_{\varepsilon}\quad\Rightarrow\quad\mathbf{F}\psi\!\in\!\mathcal{S}_{\varepsilon}.$ (47)
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{0}|\leq|\lambda_{1}|\leq...\leq|\lambda_{d-1}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
$\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\ \varphi_{\kappa}.$ (48)
The relation
$\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\,\varphi_{\kappa}\quad\Rightarrow\quad\mathbf{C}\mathbf{F}\varphi_{k}\!=\!\mathbf{F}\mathbf{C}\varphi_{k}\!=\!\lambda_{k}\ \mathbf{F}\,\varphi_{\kappa}$
Table 2: Case d=11d\!=\!11: $\varphi_{1},\,\varphi_{2},...,\varphi_{d}$ are eigenstates of $\mathbf{F}$.
$k$ $\parallel\mathbf{F}\varphi_{k}\!-\!\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!+\!\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!-\!\mathrm{i}\varphi_{k}\parallel$   $\parallel\mathbf{F}\varphi_{k}\!+\!\mathrm{i}\varphi_{k}\parallel$
@@ III.1 A spectral decomposition of the Fourier transform
and the numerical data (see Table 2) suggest that the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ are also eigenstates of $\mathbf{F}$, namely Mehta87 ; Barker
$\mathbf{F}\varphi_{n}\!=\!(-\mathrm{i})^{n}\ \varphi_{n}.$ (49)
Particularly, the discrete Fourier transform admits the spectral decomposition
$\mathbf{F}\!=\!\sum_{n=0}^{d-1}(-\mathrm{i})^{n}\ |\varphi_{n}\rangle\langle\varphi_{n}|.$ (50)
III.2 A discrete version of the uncertainty relation
A direct consequence of (43) is
$|\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi\rangle|\approx\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.$ (51)
$\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\frac{|\langle\psi|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\psi|}{2}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon},$ (52)
that is
$\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (53)
is satisfied with approximation in $\mathcal{S}_{\varepsilon}$. See the particular cases presented in the Tables V-XIII and Figure 3.
III.3 Commutation relation of discrete creation-annihilation operators
The operators
@@ III.3 Commutation relation of discrete creation-annihilation operators
can be regarded as a discrete version of the creation (also called raising) and annihilation (also called lowering) operators
@@
used in the continuous variable case. For small $\varepsilon\!>\!0$ and any $\psi\!\in\!\mathcal{S}_{\varepsilon}$, we have $\hat{\mathbf{q}},\hat{\mathbf{p}}]\psi\approx\mathrm{i}\frac{d}{2\pi}\psi\ \Rightarrow\ [\mathbf{a}^{\dagger},\mathbf{a}]\psi\approx\psi$, that is
[…]
In the continuous-variable case, by using Baker–Campbell–Hausdorff formula, one gets
$\hat{q},\hat{p}]={\rm i}\mbox{\small$\frac{h}{2\pi}$}\quad\Longrightarrow\quad D(q,p)\!=\!\mathrm{e}^{-\frac{\pi\mathrm{i}}{h}qp}\,\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}p\hat{q}}\,\mathrm{e}^{-\frac{2\pi\mathrm{i}}{h}q\hat{p}}\!=\!\mathrm{e}^{\frac{2\pi\mathrm{i}}{h}(p\hat{q}-q\hat{p})}$. (57)
By denoting $\alpha\!=\!\sqrt{\frac{\pi}{h}}(q\!+\!\mathrm{i}p),$ this relation can be written as
$D(\alpha)\!=\!\mathrm{e}^{\alpha\hat{a}^{\dagger}-\bar{\alpha}\hat{a}}$, (58)
It is an open issue if
[…]
IV Some remarkable elements of $\mathcal{S}_{\varepsilon}$ and their properties
IV.1 Discrete-variable Gaussian functions
Figure 1: Gaussian functions $\mathbf{g}_{3}$ and $\mathbf{g}_{1/3}$ in the case d=31d\!=\!31.
$\begin{array}[]{l}g_{\kappa}(q)\!=\!\mathrm{e}^{-\frac{\kappa\pi}{h}q^{2}},\end{array}$ (60)
satisfies the relation
$F[g_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\,g_{1/\kappa}.$ (61)
The corresponding discrete-variable Gaussian function $\mathfrak{g}_{\kappa}\!:\!\{-s,\!-s\!+\!\!1,...,s\!-\!\!1,\!s\}\!\rightarrow\!\mathbb{R},$
$\begin{array}[]{l}\mathfrak{g}_{\kappa}(n)=\sum\limits_{\alpha=-\infty}^{\infty}\mathrm{e}^{-\frac{\kappa\pi}{d}(n+\alpha d)^{2}},\end{array}$ (62)
satisfies the relation Vourdas04 ; Cotfas12
$\mathbf{F}[\mathfrak{g}_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\ \mathfrak{g}_{1/\kappa}$ (63)
similar to (61). It is a particular case of Jacobi $\theta$ function Vourdas04 ; Mehta87 . Because $\mathbf{F}$ is a unitary transform, the corresponding normalized discrete Gaussian functions (see Figure 1)
[…]
satisfy the relation
$\mathbf{F}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}.$ (65)
For $d$ large enough and small $\delta>0$, there exists $\eta>0$ such that, for $\kappa\!\in\!(\frac{1}{\eta},\eta)$, the most significant coordinates of $\mathbf{g}_{\kappa}$ in the eigenbasis $\{\varphi_{1},\,\varphi_{2},\,...,\,\varphi_{d}\}$ of $\mathbf{C}\!=\![\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}$ are those corresponding to the functions $\varphi_{k}$ with $|\lambda_{k}|<\delta$ (see Table 3). We can consider that
$\hat{\mathbf{q}},\hat{\mathbf{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \mathbf{g}_{\kappa}\ \mbox{with}\ \ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ (66)
(see Tables 3, 4 and Figure 2). From (49), we get
$|\langle\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\theta_{n}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\mathbf{F}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\varphi_{n}|\mathbf{g}_{1/\kappa}\rangle|,$ (67)
in agreement with the numerical data from Table 3.
@@ IV.1 Discrete-variable Gaussian functions
Since $\mathbf{F}$ is a unitary transform, we get
@@
$\parallel\mathbf{C}\mathbf{g}_{\kappa}\parallel=\parallel\mathbf{C}\mathbf{g}_{1/\kappa}\parallel,$ (69)
@@
$\langle\mathbf{g}_{1/\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{1/\kappa}\rangle\!=\!\langle\mathbf{F}\mathbf{g}_{\kappa}|\mathbf{F}[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle\!=\!\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle,$ (70)
@@
Figure 2: Norm $\parallel([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi})\mathbf{g}_{k}||$ for d=11d\!=\!11 and d=31d\!=\!31.
Table 4: $\hat{\mathbf{q}},\hat{\mathbf{p}}]\mathbf{g}_{\kappa}\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\mathbf{g}_{\kappa}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $\left([\hat{\mathbf{q}},\hat{\mathbf{p}}]-{\rm i}\frac{d}{2\pi}\right)\mathbf{g}_{\kappa}||$
@@
Table 5: $|\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle|\approx\frac{d}{2\pi}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $|\ |\langle\mathbf{g}_{\kappa}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{\kappa}\rangle|-\frac{d}{2\pi}\ |$
@@
$\hat{\mathbf{q}},\hat{\mathbf{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{and}\quad\Delta\hat{\mathbf{q}}\ \Delta\hat{\mathbf{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (71)
are satisfied with approximation for any $\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}$ (See Figure 2 and Figure 3).
Figure 3: Modulus $|\,|\langle\mathbf{g}_{k}|[\hat{\mathbf{q}},\hat{\mathbf{p}}]|\mathbf{g}_{k}\rangle|-\frac{d}{2\pi}\ |$ for d=11d\!=\!11 and d=31d\!=\!31.
@@
$\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}\quad\Rightarrow\quad\mathbf{F}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}\!\in\!\mathcal{G}_{\delta}.$ (72)
IV.2 Discrete version of Fock states